# Accurate Pocket Identification for Binding-Site-Agnostic Docking

### v1

@@ III Model Training and Inference Pipeline
Recently, PUResNet V2 34 sought to improve performance by replacing the Dice loss function with a focal loss 57. However, our extensive experiments indicate that adopting a “soft” approach 54; 55; 56 is better suited in this context. In contrast to the binary “yes/no” classification of occupied voxels, the density of pseudoatoms serves as an indicator of the proximity of a voxel to the interior or boundary of a pocket. Higher densities typically occur near polar or charged residues and ligand functional groups, whereas lower densities are often associated with hydrophobic regions.
Drawing inspiration from medical image segmentation methods 54; 55; 56, we replace the sigmoid output in the final layer with a ReLU 58 (hence the name of our model, ReLU Activated Pocket Identification for Docking, RAPID-Net):
@@
For the model training, we also replace the conventional Dice loss in Eq. 1 with its “soft” variant:
@@
Several ways of implementing the “soft” Dice loss have been proposed 55; 56, but we found that this simplest version, based on the $\text{L}^{2}$ norm, works best.
To mitigate class imbalance, as the number of non-interacting voxels sharply exceeds the number of interacting ones, we applied class reweighting using scikit-learn 59. During inference, a voxel is classified as part of a binding pocket if the model output exceeds $\texttt{threshold}\,=0.5$. When ensembling five models, we apply morphological closing using the binary_closing function from the scipy.ndimage package 60 to mitigate sparsity in the predicted pocket regions across model replicas.
Unlike previous approaches that use cavity6.mol2 labels 37; 38, we train our model with threshold-less pocket labels (cavityALL.mol2). In cavity6.mol2, annotations are restricted to regions within $6.5\,\text{\AA}$ of the ligand’s heavy atoms 49, potentially overlooking distal functional regions such as allosteric pockets, exosites, or flexible loops. These sites can critically influence drug binding and resistance, making them key therapeutic targets. Secondary binding sites, which often lie beyond the $6.5\,\text{\AA}$ boundary, mediate interactions with larger substrates or cofactors and help position them within the catalytic pocket, as further discussed in Section IX. By adopting threshold-less labels, our model identifies both catalytic and more distant sites, thereby improving predictions for proteins that contain such secondary binding regions.
This approach also makes the Distance Center Center (DCC) metric, which was previously used to evaluate Kalasanty 38 and PUResNet 37, less relevant. DCC defines the distance between the geometrical center of the predicted pocket and the true ligand binding pose, but “tunnels” can extend from the primary binding pocket to distal residues that lie far from the ligand but have significant therapeutic implications (see Section IX for examples). Furthermore, the threshold-less pocket definition is less sensitive to the ligand size, thereby improving the generalizability of our model.
IV Docking Protocol and its Rationale
@@
Fig. 5 illustrates this setup for the majority-voted pocket in the ABHD5 protein with a threshold value of $2$.
Figure 5: Majority-voted pocket predicted by RAPID-Net and its corresponding search grid with $2$ threshold for ABHD5 protein.
@@
To demonstrate that the accuracy of our model is primarily a result of architectural and training improvements–not merely ensembling–we also provide single-model predictions from Kalasanty 38, PUResNet 37, and individual RAPID-Net runs. Similarly, for each predicted pocket, we generate grids with $2\,\text{\AA}$, $5\,\text{\AA}$, $10\,\text{\AA}$, and $15\,\text{\AA}$ thresholds.
@@
V Test Benchmarks and Evaluation Metrics
@@
Docking methods like AutoDock Vina 12 balance speed and accuracy 62, utilizing “fast and dirty” scoring functions. A common strategy to enhance their accuracy is to generate an ensemble of potential poses and then identify the best pose among them using more accurate scoring tools, for example 63; 64. Fast scoring functions enable the efficient exploration of numerous poses, while more accurate methods are reserved for refinement and final pose selection. To account for this approach, we also compute the ensemble RMSD accuracy as “at least one pose in the ensemble is correct”, highlighting the potential gains from precise rescoring. Since our primary focus is on improving pocket identification rather than developing new rescoring functions, we defer such work to the future. Our goal here is to illustrate the potential of our method when combined with an appropriate rescoring tool.
Additionally, we evaluate the PB-success rate, defined as the percentage of poses with $\text{RMSD}<2$Å that pass all 26/26 chemical validity checks defined by PoseBusters 28.
Finally, similarly to 37, we compute the Pocket-Ligand Intersection (PLI) score to quantify the proportion of ligand atoms residing within the predicted pocket. Unlike 37, which computes a voxel-based intersection, we measure the average fraction of ligand heavy atoms located within $5$ of at least one pocket pseudoatom, yielding a more intuitive metric:
@@
However, unlike 37, we compute this ratio for all protein-ligand pairs in the dataset rather than restricting it to cases where the ligand is centered within $\leq 4\,\text{\AA}$ of the pocket center. If no pockets are predicted, the PLI is set to zero. When a protein has multiple predicted pockets or multiple ligand poses, we report the maximum PLI value across all possibilities. For completeness, we also indicate how many proteins have at least one predicted pocket and how many have at least one ligand pose within the defined search grids.
@@
Our first dataset to evaluate docking efficiency is the PoseBusters 28 dataset, which consists of complexes released since 2021. Since these complexes are novel and were not available when current docking programs were developed, the PoseBusters 28 dataset provides a rigorous assessment of their generalization abilities. For instance, several docking algorithms that claimed superior accuracy on CASF 2016 62 dataset, perform drastically worse on PoseBusters 28. Additionally, we evaluate docking accuracy on the Astex Diverse Set 29. Although this dataset is older than PoseBusters 28, it remains a challenging benchmark for many algorithms that had previously reported high accuracy on CASF 2016 62.
@@
We distinguish between two docking scenarios: those that use “prior knowledge” of the binding site (which is typically unavailable in real-world applications) and truly “blind” settings that lack this information. In both the PoseBusters 28 and Astex 29 datasets, when the binding site is assumed to be known, the reference ligand coordinates are used to center a $25\,\text{\AA}\times 25\,\text{\AA}\times 25\,\text{\AA}$ bounding box. However, to evaluate RAPID-Net under truly “blind”, binding-site-agnostic conditions, we omit these reference ligands entirely.
@@
In addition to docking accuracy, for completeness, we report the PLI rates for PoseBusters 28 and Astex 29 datasets. For direct comparison with PUResNet 37, we also evaluate PLI for the Coach420 30 and BU48 31 datasets, excluding any structures that appear in the training set as specified in 37. Lastly, to highlight the advantages of our model architecture and training approach, we report docking accuracy for both single RAPID-Net models and their ensembled version. PLI is calculated for individual models, majority-voted pockets, and minority-reported pockets.
VI Evaluation on the PoseBusters Dataset
@@
We obtained the following results by docking on the PoseBusters 28 dataset using the aforementioned protocol. Fig. 7 shows the percentage of predictions with $\text{RMSD}<2\,\text{\AA}$ and the corresponding PB-success rates. When guided by the ensembled RAPID-Net, AutoDock Vina 12 achieved 55.8% RMSD-correct predictions and a 54.9% PB-success rate, outperforming DiffBindFR 43, which scored 50.2% and 49.1%, respectively. This result highlights the crucial impact of accurate pocket identification on the subsequent docking accuracy. Our approach uses targeted docking with the widely used AutoDock Vina 12, whereas DiffBindFR 43 scans the entire protein surface.
@@
Figure 7: Comparison of Top-1 Vina accuracies when guided by different pocket identification algorithms, and a DiffBindFR-Smina 43, for PoseBusters 28 dataset.
@@
Furthermore, as shown in Fig. 7, docking driven by individual RAPID-Net models–the ensemble components–outperforms PUResNet 37 and Kalasanty 38 in docking accuracy. Ensembling RAPID-Net models further improves docking accuracy by providing more stable results. The distribution of RMSD values for the Top-1 predicted Vina poses is shown in Fig. 8.
Note the issue of low recall discussed above. While the RAPID-Net ensemble predicts pockets in all 308 structures, PUResNet 37 and Kalasanty 38 predict pockets in only 278 and 263 protein structures, respectively. Since we evaluate accuracy in “blind” docking settings, cases with no predicted pockets are treated as failures by default.
Not all predicted pockets are positioned correctly. Of the 308 structures where RAPID-Net predicts at least one pocket, 307 contain at least one true ligand binding pose entirely within a search grid, meaning that docking can be successful in 307 out of 308 cases. In the remaining case, docking would inevitably fail due to an inaccurate search grid location. By contrast, PUResNet 37 and Kalasanty 38 have only 265 and 249 structures, respectively, where at least one true ligand pose is located completely within the search grid. This demonstrates that our model not only accurately identifies pockets in favorable protein structures but also provides reliable predictions across the entire dataset.
@@
Figure 13: Ranks of subleading Vina poses that pass the RMSD $<$ 2 Å test when the Top-1 Vina pose fails for PoseBusters 28 dataset.
Figure 14: For the 7PRI protein structure, the Top-1 Vina pose fails while a subleading one succeeds. In (a), the majority-voted pockets are shown in cyan, and one of the true ligand binding poses largely overlaps with one of them. In (b), minimally-reported pockets are displayed.
@@
Figure 18: Distribution of search grids leading to poses that pass the RMSD $<$ 2 Å test, for Top-1 Vina poses and for subleading poses when the Top-1 Vina pose fails for PoseBusters 28 dataset.
Figure 19: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 37, and Kalasanty 38 for the PoseBusters dataset 28.
@@
Similarly to the 8BTI structure shown in Fig. 10, surprisingly many protein structures have a pattern when the Top-1 Vina pose fails the RMSD test, but one of the subleading poses passes it. As shown in Fig. 12, when evaluating RMSD as “at least one pose among the ensemble is correct”, AutoDock Vina 12 actually outperforms AlphaFold 3 33 with the pocket specified, achieving 93.8% versus 93.2% RMSD $<$ 2 Å accuracy rate. This difference becomes even more pronounced in the “blind” docking settings, where AutoDock Vina 12, guided by the ensemble of RAPID-Net models, results in a 92.2% RMSD $<$ 2 Å success rate compared to 80.5% for AlphaFold 3 33. Furthermore, RAPID-Net performs better than PUResNet 37 and Kalasanty 38 in guiding AutoDock Vina 12, as can be observed in Fig. 12.
To illustrate the extent of Vina’s limitations in pose ranking, Fig. 13 shows the distribution of ranks for subleading poses that pass the RMSD $<$ 2 Å test. In particular, poses ranked far from Top-1 sometimes succeed. For illustration, we consider the cases where the Top-1 Vina pose fails but poses ranked between 31 and 35, very far from the top, successfully pass the RMSD $<$ 2 Å test. An example of this was previously shown in Fig. 2.
@@
These results suggest that the AutoDock Vina 12 sampling mechanism is robust and reliably generates near-native binding poses within the ensemble. However, its main limitation is its ranking power: although it reliably generates poses, it struggles to distinguish correct poses from incorrect ones in its ranking.
This limitation highlights a strong potential for future improvements in our combined scheme. Integrating a more accurate reweighting or rescoring tool than the Vina scoring function could enhance the selection of the best pose, leading to improved overall accuracy. This approach offers a promising alternative to computationally expensive docking tools such as AlphaFold 3 33, potentially achieving similar or better performance with less computational overhead.
The distributions of the grids leading to RMSD-correct Top-1 poses and subleading poses (when the Top-1 pose fails) are shown in Fig. 18. Notably, a large portion of the larger grids results in poses that pass the RMSD $<$ 2 Å test.
Finally, Fig. 19 shows the distribution of PLI values for PUResNet 37, Kalasanty 38, and RAPID-Net for this dataset. All five replicas of RAPID-Net outperform both PUResNet 37 and Kalasanty 38. Majority voting stabilizes the PLI predictions, and including minority-reported pockets further improves the PLI.
@@
In the next Section VII, we perform docking on the Astex Diverse Set 29 following a similar procedure as described for the PoseBusters 28 dataset.
VII Evaluation on the Astex Diverse Set
Figure 20: Comparison of docking accuracies for the Astex Diverse Set 29 when AutoDock Vina 12 is guided by different pocket prediction algorithms.

### v2

@@ IV Model Training and Inference Pipeline
Recently, PUResNet V2 66 sought to improve performance by replacing the Dice loss function with a focal loss 89. However, our extensive experiments indicate that adopting a “soft” approach 86; 87; 88 is better suited in this context. In contrast to the binary “yes/no” classification of occupied voxels, the density of pseudoatoms serves as an indicator of the proximity of a voxel to the interior or boundary of a pocket. Higher densities typically occur near polar or charged residues and ligand functional groups, whereas lower densities are often associated with hydrophobic regions.
Drawing inspiration from medical image segmentation methods 86; 87; 88, we replace the sigmoid output in the final layer with a ReLU 90 (hence the name of our model, ReLU Activated Pocket Identification for Docking, RAPID-Net):
@@
For the model training, we also replace the conventional Dice loss in Eqn. 1 with its “soft” variant:
@@
Several ways of implementing the “soft” Dice loss have been proposed 87; 88, but we found that this simplest version, based on the $\text{L}^{2}$ norm, works best.
To mitigate class imbalance, as the number of non-interacting voxels sharply exceeds the number of interacting ones, we applied class reweighting using scikit-learn 91. During inference, a voxel is classified as part of a binding pocket if the model output exceeds $\texttt{threshold}\,=0.5$. When ensembling five models, we apply morphological closing using the binary_closing function from the scipy.ndimage package 92 to mitigate sparsity in the predicted pocket regions across model replicas.
Unlike previous approaches that use cavity6.mol2 labels 69; 70, we train our model with threshold-less pocket labels (cavityALL.mol2). In cavity6.mol2, annotations are restricted to regions within $6.5\,\text{\AA}$ of the ligand’s heavy atoms 81, potentially overlooking distal functional regions such as allosteric pockets, exosites, or flexible loops. These sites can critically influence drug binding and resistance, making them key therapeutic targets. Secondary binding sites, which often lie beyond the $6.5\,\text{\AA}$ boundary, mediate interactions with larger substrates or cofactors and help position them within the catalytic pocket, as further discussed in Section X. By adopting threshold-less labels, our model identifies both catalytic and more distant sites, thereby improving predictions for proteins that contain such secondary binding regions.
This approach also makes the Distance Center Center (DCC) metric, which was previously used to evaluate Kalasanty 70 and PUResNet 69, less relevant. DCC defines the distance between the geometrical center of the predicted pocket and the true ligand binding pose, but “tunnels” can extend from the primary binding pocket to distal residues that lie far from the ligand but have significant therapeutic implications (see Section X for examples). Furthermore, the threshold-less pocket definition is less sensitive to the ligand size, thereby improving the generalizability of our model.
V Pocket-informed Docking Protocol and its Rationale
@@
Fig. 4 illustrates this setup for the majority-voted pocket in the ABHD5 protein 93 with a threshold value of $2$.
Figure 4: Majority-voted pocket predicted by RAPID-Net and its corresponding search grid with $2$ threshold for ABHD5 protein 93.
@@
To demonstrate that the accuracy of our model is primarily a result of architectural and training improvements–not merely ensembling–we also provide single-model predictions from Kalasanty 70, PUResNet 69, and individual RAPID-Net runs. Similarly, for each predicted pocket, we generate grids with $2\,\text{\AA}$, $5\,\text{\AA}$, $10\,\text{\AA}$, and $15\,\text{\AA}$ thresholds.
@@
VI Test Benchmarks and Evaluation Metrics
@@
Docking methods like AutoDock Vina 12 balance speed and accuracy 95, utilizing “fast and dirty” scoring functions. A common strategy to enhance their accuracy is to generate an ensemble of potential poses and then identify the best pose among them using more accurate scoring tools, for example 96; 97. Fast scoring functions enable the efficient exploration of numerous poses, while more accurate methods are reserved for refinement and final pose selection. To account for this approach, we also report the sampling RMSD accuracy, defined as the percentage of cases with “at least one correct pose in the ensemble”. This metric underscores the potential gains that can be achieved through more accurate rescoring. Since our primary focus is on improving pocket identification rather than developing new rescoring functions, we defer such work to the future. Additionally, we evaluate the PB-success rate, defined as the percentage of poses with $\text{RMSD}<2$ that pass all chemical validity checks defined by PoseBusters 28.
Finally, similarly to 69, we compute the Pocket-Ligand Intersection (PLI) score to quantify the proportion of ligand atoms residing within the predicted pocket. Unlike 69, which computes a voxel-based intersection, we measure the average fraction of ligand heavy atoms located within $5$ of at least one pocket pseudoatom, yielding a more intuitive metric:
@@
However, unlike 69, we compute this ratio for all protein-ligand pairs in the dataset rather than restricting it to cases where the ligand is centered within $\leq 4\,\text{\AA}$ of the pocket center. If no pockets are predicted, the PLI is set to zero. When a protein has multiple predicted pockets or multiple ligand poses, we report the maximum PLI value across all possibilities. For completeness, we also indicate how many proteins have at least one predicted pocket and how many have at least one ligand pose within the defined search grids.
@@
Our primary dataset for evaluating docking performance is the PoseBusters 28 dataset, which consists of novel protein-ligand complexes that were not available during the development of current docking programs and thus provides a rigorous assessment of their generalization capabilities. For instance, several docking algorithms that claimed superior accuracy on CASF 2016 95 dataset, perform drastically worse on PoseBusters 28. Additionally, we evaluate docking accuracy on the Astex Diverse Set 33. Although this dataset is older than PoseBusters 28, it remains a challenging benchmark for many algorithms that had previously reported high accuracy on CASF 2016 95.
@@
We distinguish between two docking scenarios: those that use “prior knowledge” of the binding site (which is typically unavailable in real-world applications) and truly “blind” settings that lack this information. In both the PoseBusters 28 and Astex 33 datasets, when the binding site is assumed to be known, the reference ligand coordinates are used to center a $25\,\text{\AA}\times 25\,\text{\AA}\times 25\,\text{\AA}$ bounding box. However, to evaluate RAPID-Net under truly “blind”, binding-site-agnostic conditions, we omit these reference ligands entirely.
@@
In addition to docking accuracy, for completeness, we report the PLI rates for PoseBusters 28 and Astex 33 datasets. For direct comparison with PUResNet 69, we also evaluate PLI for the Coach420 34 and BU48 35 datasets, excluding any structures that appear in the training set as specified in 69. Lastly, to highlight the advantages of our model architecture and training approach, we report docking accuracy for both single RAPID-Net models and their ensembled version. PLI is calculated for individual models, majority-voted pockets, and minority-reported pockets.
VII Evaluation on the PoseBusters Dataset
@@
We obtained the following results by docking the PoseBusters 28 dataset using the aforementioned protocol. Fig. 6 shows the percentage of predictions with $\text{RMSD}<2\,\text{\AA}$ and the corresponding PB-success rates. When guided by the ensembled RAPID-Net, AutoDock Vina 12 achieved 55.8% RMSD-correct predictions and a 54.9% PB-success rate, outperforming DiffBindFR 27, which scored 50.2% and 49.1%, respectively. This result highlights the crucial impact of accurate pocket identification on the subsequent docking accuracy. Our approach uses targeted docking with the widely used AutoDock Vina 12, whereas DiffBindFR 27 scans the entire protein surface.
@@
Figure 6: Comparison of Top-1 Vina accuracies when guided by different pocket identification algorithms, and a DiffBindFR-Smina 27, for PoseBusters 28 dataset.
@@
Figure 8: Performance comparison on the subset of novel PoseBusters structures as suggested in 31. The RAPID-Net-guided AutoDock Vina achieves 53.1% compared to 59.5% of AlphaFold 3, with significantly lower computational costs.
@@
Furthermore, as shown in Fig. 6, docking driven by individual RAPID-Net models–the ensemble components–outperforms PUResNet 69 and Kalasanty 70 in docking accuracy. Ensembling RAPID-Net models further improves docking accuracy by providing more stable results. The distribution of RMSD values for the Top-1 predicted Vina poses is shown in Fig. 7. Note the issue of low recall discussed above. While the RAPID-Net ensemble predicts pockets in all 308 structures, PUResNet 69 and Kalasanty 70 predict pockets in only 278 and 263 protein structures, respectively. Since we evaluate accuracy in “blind” docking settings, cases with no predicted pockets are treated as failures by default.
Not all predicted pockets are positioned correctly. Of the 308 structures where RAPID-Net predicts at least one pocket, 307 contain at least one true ligand binding pose entirely within a search grid, meaning that docking can be successful in 307 out of 308 cases. In the remaining case, docking would inevitably fail due to an inaccurate search grid location. By contrast, PUResNet 69 and Kalasanty 70 have only 265 and 249 structures, respectively, where at least one true ligand pose is located completely within the search grid. This demonstrates that our model not only accurately identifies pockets in favorable protein structures but also provides reliable predictions across the entire dataset.
For comparison, Fig. 8 presents the evaluation on a subset of novel protein-ligand complexes from the PoseBusters benchmark, as defined by31. Compared to the best-performing structure prediction model, AlphaFold 3, as evaluated by PoseBench, RAPID‑Net + Vina achieves a Top-1 success rate of 53.1%, closely approaching AF3’s 59.5%. This result highlights RAPID-Net’s ability to achieve competitive accuracy while offering a substantially lower computational cost.
@@
Figure 13: For the 7PRI protein structure, the Top-1 Vina pose fails while a subleading one succeeds. In (a), the majority-voted pockets are shown in cyan, and one of the true ligand binding poses largely overlaps with one of them. In (b), minimally-reported pockets are displayed as purple sticks.
@@
Figure 17: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 69, and Kalasanty 70 for the PoseBusters dataset 28.
@@
Similarly to the 8BTI example in Fig. 10, many protein structures have the following pattern: the Top–1 AutoDock Vina pose fails the RMSD $<$ 2 Å criterion, yet a lower-ranked pose meets it. To illustrate this behavior, we report the sampling accuracy, defined as the fraction of cases in which at least one pose in the ensemble achieves RMSD $<2\,\text{\AA}$. Under this metric, as shown in Fig. 12, AutoDock Vina with the prior knowledge of the search box 12 has a success rate of 93.8%.
In comparison, AlphaFold 3, when run with pocket information available, achieves a Top–1 accuracy of 93.2% with RMSD $<2\,\text{\AA}$ 30. We did not compute AlphaFold 3’s sampling accuracy and therefore make no claim of overall superiority. This comparison merely illustrates that the current limitations in docking accuracy stem more from pose ranking than from conformational sampling.
This difference widens in fully “blind” docking settings: AutoDock Vina, guided by a RAPID–Net, achieves the 92.2% of RMSD $<2\,\text{\AA}$ sampling accuracy, whereas AlphaFold 3 achieves 80.5% of RMSD $<2\,\text{\AA}$ Top–1 accuracy. These results suggest that improved re-ranking algorithms could further boost pocket-guided docking workflows. Furthermore, RAPID-Net performs better than PUResNet 69 and Kalasanty 70 in guiding AutoDock Vina 12, as can be observed in Fig. 12.
To better illustrate this pattern, we consider four representative cases in which the Top–1 AutoDock Vina pose fails to reach the RMSD $<2\,\text{\AA}$ threshold, while a pose with a much lower ranking succeeds.
@@
These results illustrate that the AutoDock Vina 12 sampling mechanism is robust and reliably generates near-native binding poses within the ensemble. However, its main limitation is its ranking power: although it reliably generates poses, it struggles to distinguish correct poses from incorrect ones in its ranking.
Finally, Fig. 17 shows the distribution of PLI values for PUResNet 69, Kalasanty 70, and RAPID-Net for this dataset. All five replicas of RAPID-Net outperform both PUResNet 69 and Kalasanty 70. Majority voting stabilizes the PLI predictions, and including minority-reported pockets further improves the PLI.
@@
In the next Section VIII, we perform docking on the Astex Diverse Set 33 following a similar procedure as described for the PoseBusters 28 dataset.
VIII Evaluation on the Astex Diverse Set
Figure 18: Comparison of docking accuracies for the Astex Diverse Set 33 when AutoDock Vina 12 is guided by different pocket prediction algorithms.